OpenStudy (anonymous):

f(x) = 2x^4 – x^2 – 1 By using the substitution u = x^2, find the horizontal intercepts.

OpenStudy (bahrom7893):

horizontal intercepts mean y = 0: 0 = 2x^4 - x^2 -1 using substitution u = x^2, we get 0 = 2u^2 - u - 1, solve this equation and post your steps.

OpenStudy (dan815):